OpenStudy (anonymous):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)?

OpenStudy (mathmale):

I'll share some background info with you here: 1. The standard form of a circle with center at (0,0) and radius r is\[x^2+y^2=r^2\] 2. The std. form of a circle with center at (h,k) and radius r is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (mathmale):

If you'd substitute the given info into the 2nd equation properly, you'll be able to calculate the radius, r. Then write the equation of this circle including the now-known value of r.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (mathmale):

Please let me know if and when you need further help in solving this problem.